1. Consider the following demand and supply functions for vitamins :

Qd= 100 - 5P and Qs= 4 + 3P.

  1. Graph the supply and demand functions in the typical manner with price (P) on the Y-axis and quantity on the X-axis, showing their intercepts.
  2. What is the slope of each line?
  3. What is the equilibrium price and quantity?
Answer #1

Answer

a)

to graph it

we need to plot two-point of each line and then join the line

for demand

P=0

Q=100-5*0=100

Q=0 then 0=100-5P

5P=100

P=20

=======

for supply

P=0

Q=4+3*0=4

P=20

Q=4+3*20=64

=====

b)

The slope of the line is found by differentiation

The slope of demand is -1/5 and supply is 1/3

=====

c)

equilibrium is at Qd=Qs

100-5P=4+3P

8P=96

P=12

Q=4+3*12=40

the equilibrium quantity is 40 units and price is 12